Hess’s Law (a) Calculate the ∆H for the reaction: CCl4(g) → C(s, graphite) + 2Cl2(g)

using the following chemical equations and their respective enthalpy changes:

C(s, graphite) + 2F2(g) → CF4(g) ∆H = -679.9 kJ

CF4(g) + 2Cl2(g) → CCl4(g) + 2F2(g) ∆H = 573.2 kJ

(b) Calculate the ∆H for the reaction: C(s, graphite) + 2H2O(g) → CH4(g) + O2(g)

using the following chemical equations and their respective enthalpy changes:

C(s, graphite) + O2(g) → CO2(g) ∆H = -394 kJ

CH4(g) + 2O2(g) → CO2(g) + 2H2O(g) ∆H = -802 kJ
Answer #1

a) Let; we have given;

C(s, graphite) + 2F2(g) CF4(g)H =-679.9kJ ------(1)

CF4(g) + 2Cl2(g)CCl4(g) + 2F2(g)   H = 573.2kJ ----(2)

Reversing (1) and (2) and adding we get;

CF4(g) C(s, graphite) + 2F2(g)H =+679.9kJ ------(3) new reaction

CCl4(g) + 2F2(g)CF4(g) + 2Cl2(g)H = -573.2kJ ----(4)​ new reaction

------------------------------------------------------------------

CCl4(g) C(s, graphite) + 2Cl2(g) H =+106.7kJ​

b) Let; we have given;

C(s, graphite) + O2(g) CO2(g)H =-394kJ ------(1)

CH4(g) + 2O2(g) CO2(g) + 2H2O(g)   H = -802kJ ----(2)

Reversing (2) and adding we get;

C(s, graphite) + O2(g) CO2(g)H =-394kJ ------(1)

CO2(g) + 2H2O(g) CH4(g) + 2O2(g)H = +802 kJ ----(3)​ new reaction

------------------------------------------------------------------

C(s, graphite) + 2H2O(g) CH4(g) + O2(g) H =+408kJ​